OpenStudy (anonymous):

what is the transformation of the graph g(x)=-log(x)+2

OpenStudy (debbieg):

Do you know what y=log(x) looks like? If so, then what are your transformations here?

OpenStudy (anonymous):

i dont understand

OpenStudy (debbieg):

First go from y=log(x) ---to---> y=-log(x) What does that do to the graph? Then from y=-log(x) ---to---> y=-log(x)+2 What does that do to the graph?

OpenStudy (anonymous):

it adds 2 to the right

OpenStudy (anonymous):

right?

OpenStudy (debbieg):

no.... but start with the first one. y=log(x) ---to---> y=-log(x) What does that do to the graph? If I have y=f(x) and then I go to y=-f(x), what do I do? First I evaluate f(x) (just like before the transformation), then I take every y value and multiply it by (-1).

OpenStudy (debbieg):

What does that do to the graph?

OpenStudy (anonymous):

flipped it

OpenStudy (debbieg):



OpenStudy (debbieg):

Right, flips it over the x-axis.

OpenStudy (anonymous):

ok i get that part now

OpenStudy (debbieg):

Now I take f(x) and add 2 to it... what does that do...? What happens to every y-coordinate?

OpenStudy (anonymous):

2 gets added

OpenStudy (debbieg):

Yes, 2 gets added to the y coordinates.... so the graph does what?

OpenStudy (debbieg):

does it stretch, shrink, shift up, down, shift right, left? .... what?

OpenStudy (anonymous):

flips over the x- axis and moves two to the right